OpenStudy (anonymous):

Solve the triangle. A = 32°, a = 19, b = 12

OpenStudy (johnweldon1993):

Remember the law of sines....a/A = b/B = c/C so if 32/19 is A then 32/19 = x/12 after that...you'll have 2 of the 3 angles...with that...you know that the angles of a triangle add up to what?

OpenStudy (anonymous):

180 Degrees

OpenStudy (johnweldon1993):

correct...so now...for A to correct myself as well side A divided by the sin of angle A means what?

OpenStudy (anonymous):

I'm not sure..

OpenStudy (johnweldon1993):

the law of sines....means you take... \[a/\sin A=b/\sin B=c/\sin C\]

OpenStudy (johnweldon1993):

where a, b, and c are side lengths...and A, B, and C are angles

OpenStudy (johnweldon1993):

so for the first part.... a/sinA....what is your a and what is your A?

OpenStudy (anonymous):

A=32 a=19

OpenStudy (johnweldon1993):

correct...so now to solve that...you calculate..... \[19/\sin 32\]

OpenStudy (anonymous):

so thats about35.85

OpenStudy (johnweldon1993):

correct...so now for the second part...we know what b is...but not what B is....so in the equation...b/sinB....it is 12/sinB remember how you just got 35.85? 35.85 is a/sinA right? looking at the law of sines now...we have 35.85 = 12/sinB how do we solve for sinB?

OpenStudy (anonymous):

I also found the Sin of B to be 2.9 by setting the 19/sin32=12/sinB

OpenStudy (johnweldon1993):

hmmm lets look at your calculation for a second.... \[19/\sin32 = 12/sinB\] to solve for sinB...we multiply both sides of the equation by sinB...then dividie both sides by 19/sin32 to isolate sinB so we have sinB = 12/19sin32

OpenStudy (johnweldon1993):

but do we WANT sinB? no we want B.....so how do we isolate B? we take the inverse sin of whatever answer you get for 12/19sin32

OpenStudy (anonymous):

Im not sure how to make the inverse of 0.3483

OpenStudy (johnweldon1993):

....do you have a calculator on you? if so..just do the second button and press sin...most likely the same for many calculators

OpenStudy (anonymous):

oh, i apologize, i could use the arcsine, i just did'nt think about that one. sorry

OpenStudy (johnweldon1993):

no problem :) okay so what do you get from that?

OpenStudy (anonymous):

I found that to be .3557

OpenStudy (johnweldon1993):

make sure your calculator is in degree mode and not radian mode

OpenStudy (anonymous):

got it! 20.382

OpenStudy (johnweldon1993):

there you go :) okay....so thats your angle B correct?

OpenStudy (anonymous):

yes!. so from there I can find the others using the same method?

OpenStudy (johnweldon1993):

precisely....remembering to find the third angle....180 - the 2 you already found

OpenStudy (anonymous):

got it. thanks so much.

OpenStudy (johnweldon1993):

anytime :) let me know what you end up with

OpenStudy (anonymous):

B=20.38, C=127.6, c=27.3

OpenStudy (johnweldon1993):

with the rounding you were using...you are correct :)